How can you find the equation of the axis of symmetry for this parabola
y=xsquared - 6x + 7

Answer by scott8148(6628)

for y=ax^2+bx+c, the equation of the axis of symmetry is x=-b/2a ... in your case, x=-(-6)/2(1) or x=3